### v1

# Cosmic Ray Diffusion in the Turbulent Interstellar Medium: Effects of Mirror Diffusion and Pitch Angle Scattering

Abstract
Cosmic rays (CRs) interact with turbulent magnetic fields in the intestellar medium, generating nonthermal emission. After many decades of studies, the theoretical understanding of their diffusion in the ISM continues to pose a challenge. This study numerically explores a recent prediction termed ”mirror diffusion” and its synergy with traditional diffusion mechanism based on gyroresonant scattering. Our study combines 3D MHD simulations of star-forming regions with test particle simulations to analyze CR diffusion. We demonstrate the significance of mirror diffusion in CR diffusion parallel to the magnetic field, when the mirroring condition is satisfied. Our results support the theoretical expectation that the resulting particle propagation arising from mirror diffusion in combination with much faster diffusion induced by gyroresonant scattering resembles a Levy-flight-like propagation. Our study highlights the necessity to reevaluate the diffusion coefficients traditionally adopeted in the ISM based on gyroresonant scattering alone. For instance, our simulations imply a diffusion coefficient $\sim 10^{27}cm^{2}/s$ for particles with a few hundred TeV within regions spanning a few parsecs around the source. This estimate is in agreement with gamma-ray observations, which shows the relevance of our results for understanding of diffuse gamma-ray emission in star-forming regions.
Interstellar medium(847) — Cosmic rays(329) — Particle astrophysics(96) — Magnetic fields(994) — Magnetohydrodynamics(1964)
1 Introduction
A comprehensive understanding of cosmic ray diffusion is pivotal for addressing a multitude of issues, spanning from the generation of non-thermal radiation to the analysis of chemical composition in diverse astrophysical environments, including our galaxy.
Cosmic rays (CRs) undergo the interactions with magnetic fluctuations and change the direction of their trajectories as they diffuse through the magnetized interstellar medium (ISM). The amount of time these particles spend during their journey through the ISM plays a crucial role in elucidating phenomena such as propagation, acceleration, and non-thermal emission. The commonly referred approach to studying CR diffusion is the quasi-linear theory (QLT Jokipii, 1966; Schlickeiser, 2002), which evaluates the diffusion of CR considering small perturbations in the average magnetic field.
Accounting for the actual properties of MHD turbulence (see Beresnyak & Lazarian, 2019) is essential for understanding the pitch-angle scattering. For instance, Yan & Lazarian (2002, 2004) demonstrated that fast modes are more important compared to the other modes, i.e., Alfvén and slow modes, for gyroresonant scattering of CRs by turbulence injected at the scale height of the spiral galaxies. This happens in the ionized medium in most of galactic environments, even though the fast modes are subject to collisionless damping. The variations of magnetic field strength, arising, for instance, from the large scale slow modes, induce the modification of QLT and change the scattering rate (Yan & Lazarian, 2008; Lynn et al., 2012; Xu & Lazarian, 2018).
However, QLT and its modifications still face difficulties in reproducing the observation-required diffusion coefficients (Shalchi et al., 2004; Lazarian & Beresnyak, 2006; Yan & Lazarian, 2008; Shalchi et al., 2009; Cohet & Marcowith, 2016; Krumholz et al., 2020). The puzzling results include, for instance, the analysis of H.E.S.S data. Huang et al. (2018) suggests that in the vicinity of the Vela X pulsar wind nebula, CR electrons are likely to exhibit a diffusion coefficient of $\lesssim 10^{28}cm^{2}/s$ at $10$ TeV within the inner region spanning a few tens of parsecs around the nebula. Similar estimations have been derived from HAWC observations concerning Geminga and PSR B0656+14 pulsars wind nebulae (Abeysekara et al., 2017). Even more intriguing is the recent detection by LHAASO of very high-energy (VHE) emission at hundreds of TeVs, compatible with the presence of several PeV (Pevatron) sources of CRs inside our galaxy (Cao et al., 2021). The potential sources linked to these observations include molecular cloud-supernova remnant (MC-SNR) interaction regions (e.g. Neronov et al., 2012; Blasi et al., 2012; Ackermann et al., 2013), pulsar wind nebula halos (also called PWN TeV halos) (Recchia et al., 2021; Yan et al., 2023, 2024), and Young Star Clusters (YSCs) (Morlino et al., 2021; Abeysekara et al., 2021). The Cygnus cocoon, a very extended region surrounding a cluster of massive stars (Abeysekara et al., 2021), is perhaps one of the most compelling examples of this class. Its VHE emission originates from an expansive area surrounding the stars, likely due to diffusion away from them and into the ISM. Strong assumptions are often made to explain these observations. These assumptions go from almost free parametrization of the diffusion coefficient, as being some fraction of the average ISM diffusion coefficient estimated from QLT, up to the consideration of self-confinement theories (e.g., Olmi & Bucciantini, 2019; Giacinti & Semikoz, 2023).
A long-standing theoretical difficulty with scattering is the $90^{\circ}$ problem (Jokipii, 1966). QLT falls short in accounting for the scattering when the particle’s pitch angle approaches $90^{\circ}$. Earlier investigations have explored phenomenological models of turbulence in conjunction with QLT (see e.g. Matthaeus et al., 1990; Giacalone & Jokipii, 1999; Shalchi, 2016). However, these models have faced challenges from observations and simulations (Qin et al., 2002; Gabici et al., 2019). Inevitably, the solutions ought to rely on thoroughly validated models of MHD turbulence. Such models can significantly alter pitch angle scattering, acceleration, and spatial diffusion of CRs compared to previous ad hoc approaches (e.g. Yan & Lazarian, 2002, 2004; Brunetti & Lazarian, 2007; Beresnyak et al., 2011; Xu & Yan, 2013; Cohet & Marcowith, 2016; Medina-Torrejón et al., 2021; Kadowaki et al., 2021; Fornieri et al., 2021; Kempski et al., 2023; Lemoine, 2023; Medina-Torrejón et al., 2023; Gao & Zhang, 2024).
The super-diffusive behavior of turbulent magnetic field lines (Lazarian & Vishniac, 1999; Xu & Yan, 2013; Lazarian & Yan, 2014; Beresnyak, 2013) and particle interaction with magnetic mirrors (Kulsrud & Pearce, 1969; Xu & Lazarian, 2020) have been combined in Lazarian & Xu (2021) (henceforth LX21) to predict a new diffusion mechanism which was termed “mirror diffusion”. LX21 demonstrated that in the presence of superdiffusion of magnetic field lines, the probability of CRs getting trapped and bouncing back and forth between the same magnetic mirrors is negligibly small. CRs do not retrace their trajectories and bounce every time from different mirrors. As the mirroring causes particles to reverse their moving directions, the mirror diffusion naturally solves the 90 degree problem and is a much slower diffusion process compared to the diffusion induced by gyroresonant scattering alone.
Naturally, this new process affects only the particles with Larmor radius smaller than the size of the magnetic mirrors and a pitch angle larger than the critical angle $\alpha_{pitch,c}$. This is equivalent to say that $\mu=\cos\alpha_{pitch}$ is less than some critical value which depends on the amplitude of magnetic compressions and the efficiency of pitch-angle scattering (LX21).
The magnetic compressions induced by the cascade of fast and slow modes form a hierarchy of amplitudes from large compressions $\delta B_{L}$ at the injection scale $L$ to the smallest ones $\delta B_{d}$ at the dissipation scale $l_{d}$. These magnetic compressions induce the hierarchy of magnetic mirrors of different sizes. CRs with the largest $\mu=\mu_{c}$ bounce from large-scale magnetic compressions, while CRs with smaller $\mu$ bounce from magnetic compressions at smaller scales. As mirror diffusion preserves the first adiabatic invariant $(1-mv_{\bot}^{2})/B$ (Kulsrud & Pearce, 1969), the initial value of $\mu$ does not changed, and thus, the diffusion coefficient for mirror diffusion alone is a function of $\mu$. Effective mirroring requires that the CR gyroradius is smaller than the perpendicular extension of the compressible magnetic fluctuation. As $\mu_{c}$ where the transition between mirror diffusion and scattering diffusion takes place depends on CR energy (LX21), thus, the overall CR diffusion depends on its Larmor radius, i.e., on its energy.
In highly compressible environments like molecular clouds, we expect fast modes to dominate the mirroring effect (Xu & Lazarian, 2020; Xu, 2021). The resulting diffusion coefficient for mirror diffusion shows a strong $\mu$-dependence (LX21 and Zhang & Xu, 2023). The dependence on $\mu$, makes the mirror diffusion very different from the classical resonance scattering diffusion that isotropizes $\mu$. If we are interested in the total CR diffusion flux, we need information on the distribution of $\mu$, a new element for the CR propagation paradigm. It is reasonable to assume that CRs have the isotropic $\mu$-distribution near the source. With this assumption, considering mirror diffusion, Xu (2021) estimated the spectrum of CRs in molecular clouds (MCs) with highly compressible turbulence subject to interactions with supernova remnants (SNR) and found that it leads to a steeper high-energy gamma-ray spectrum which can naturally explain observed systems like IC 443 and W44.
Certain analytical predictions from LX21 were also validated successfully in Zhang & Xu (2023). However, the crucial importance of magnetic bouncing and mirror diffusion motivates us for more detailed studies of the processes. In this work, we investigate the mirror diffusion of charged particles in a highly compressible environment using 3D-MHD simulations of an MC-like environment as the background and test particle propagation models to evaluate the diffusion behavior of the particles aiming at testing numerically the aforementioned predictions of Xu & Lazarian (2020) and LX21.
This paper is structured as follows. In Section 2, we describe our model and the setup used in our simulations, and in Section 3 we present the results of our test particle simulations. In Section 4 we discuss our results and compare them with theoretical predictions and observations, and in Section 5, we draw our conclusions.
2 Simulation setup
2.1 Background MHD turbulence
In order to obtain a fully developed turbulent magnetic field required to study diffusion of test particles, we employ a 3D-MHD model of a molecular cloud produced in Barreto-Mota et al. (2021). This model has a sonic Mach number $M_{s}=7.0$, an Alfvénic Mach number $M_{A}=0.6$ with a mean magnetic field parallel to the z direction, and a cubic box domain with periodic boundaries and a resolution of $512^{3}$. The model is isothermal and includes self-gravity in the later stages, but in this study, we consider an earlier snapshot of the simulation before self-gravity is turned-on, but with turbulence actively present and the cascade fully developed. Turbulence is driven solenoidally in the Fourier space, around a characteristic wavelength that defines the injection scale. In this particular model, the injection scale is set to $1/4L_{0}$, where $L_{0}$ corresponds to the box size.
Figure 1 shows the energy spectrum of the turbulent magnetic field for the background MHD model, after turbulence has reached a nearly steady state (see Barreto-Mota et al., 2021, for details). The blue line shows the results for the complete background magnetic field while the orange one shows the same model with a smoothed background magnetic field, where the smaller scale perturbations have been suppressed. To smooth the small scale perturbations, we calculated the 3D Fourier transform of each component of the magnetic field, set to zero the values corresponding to $k>5$, and then transformed it back to physical space. The green dashed line shows the Kolmogorov scaling with $k^{-5/3}$ for comparison. We see that a substantial part of the inertial range of the magnetic field in this smoothed model is eliminated. We will employ both models for the CR propagation analysis. The smoothing of the magnetic field reduces the effects of scattering on the propagation of particles, allowing us to focus on the effects of mirroring diffusion.
Figure 1: Magnetic energy spectrum of the magnetic field components used in our models. The blue line shows the original complete magnetic energy spectrum (Complete models as referred in the text). The orange line shows that for the smoothed magnetic field (see text for details).
2.2 Test Particle Simulations
To study the CR diffusion in the turbulent magnetic field generated in the aforementioned model, we employ the test particle code developed by Beresnyak et al. (2011) (see also Xu & Yan, 2013; Hu et al., 2022). The particles are allowed to diffuse through the nearly steady state background turbulent magnetic field of the simulated model described in Section 2.1.
To follow the propagation of the particles through the turbulent magnetic fields and derive their trajectories, we compute the Lorentz force in each position along the particle trajectory:
$\frac{d\mathbf{v}}{dt}=\frac{q}{\gamma mc}\mathbf{v}\times\mathbf{B},$ (1)
where $\mathbf{v}$ is the particle velocity, $\gamma$ its Lorentz factor $\gamma=1/\sqrt{1-v^{2}/c^{2}}$, and $\mathbf{B}$ is the magnetic field.
Each charged particle has:
$R_{L}$ $=$ $\frac{v_{\perp}}{\Omega}=\frac{\gamma mcv_{\perp}}{qB}$ (2)
where $R_{L}$ is its Larmor radius, $m$ its mass, $q$ its charge, and $\Omega$ its gyrofrequence.In this study, we consider the particle velocity approximately given by the light speed $c$. We take particles with constant Larmor radius and initial $v_{\perp,0}\sim c$.
The code uses a Burlisch-Stoer method to trace the trajectories of test particles and uses an adaptive time-step. At each time step, the magnetic field at the grid points is interpolated to the test particles position using a cubic spline routine (Press et al., 1986; Xu & Yan, 2013).
Given that the MHD turbulent models are scale free and the energy of the particles depend on the properties of the simulated box (see Section 4), one can scale the system knowing the size of the domain $L_{0}$, the mean density and the sound speed as well as the scale independent parameters $M_{A}$ and $M_{s}$ that are defined at the injection scale of turbulence.
2.3 Initial setup
The test particle models we simulated are summarized in Table 1.
Table 1: Initial conditions for all models.
Models $\mu_{0}=\cos\theta_{0}$ $R_{L}/L_{0}$ $k_{cut}$
r03mu20 0.2 0.03
r03mu80 0.8 0.03
r06mu20 0.2 0.06
r06mu80 0.8 0.06
r10mu20 0.2 0.10
r10mu80 0.8 0.10
r03mu20k5 0.2 0.03 5
r03mu80k5 0.8 0.03 5
r06mu20k5 0.2 0.06 5
r06mu80k5 0.8 0.06 5
r10mu20k5 0.2 0.10 5
r10mu80k5 0.8 0.10 5
Each model is characterized by two free parameters, namely, the initial pitch angle ($\theta_{0}$) between the particles velocity and the local magnetic field, which we consider to be either $\mu_{0}=\cos\theta_{0}=0.2$ or $0.8$, and the Larmor radius ($R_{L}$), $R_{L}/L_{0}=0.03,0.06$ and $0.10$, in terms of the number of cells of the background simulation, $R_{L}(\#cells)=15,30$, and $50$, respectively. We note that particles are not accelerated and their energy stays the same in our simulations, we focus only in their diffusion. Each model name is labeled with these two parameters in Table 1. As described in section 2.1, we also consider two sets for the magnetic field configuration. One with the entire background turbulent magnetic field distribution, and another where the small-scale perturbations of the field were suppressed for $k_{cut}=5$, which corresponds to a dissipative scale of about $\sim 0.2L_{0}$ ($\sim 100$ grid cells) [^1] Values of $k_{cut}>5$ were tested, but the results showed little to no notable differences with regard to the model with the complete distribution of the magnetic field.. These latter models are labeled $k5$ in Table 1. The dissipation scale of the complete distribution of magnetic field is around $k\sim 250$, which corresponds to $0.004L_{0}$ approximately ($\lesssim 3$ grid cells). Models with the first background are initialized with 1000 particles, while the ones with the second background are initialized with 3000 particles. This difference will be discussed in the next sections.
3 Results
3.1 Particle trajectory and mirroring identification
Figure 2: Example of a test particle trajectory (represented by solid colored circles) with $R_{L}/L_{0}\simeq 0.03$ and $\mu_{0}=0.2$ in the background with complete distribution of magnetic field. Initially, the particle travels upwards parallel to the magnetic field (black circles), reentering the domain from the bottom face until it encounters magnetic mirrors that trap it for a while (dark purple to orange circles), diffusing perpendicularly to the field lines until it escapes again traveling downwards and reentering through the upper region of the domain, finding another mirror at the end of its trajectory (bright yellow circles). A few field lines are shown in red, white, and blue for visual distinction only. The color scales on the right hand side show the time elapsed in the particle trajectory between 850 and 1450 (same units as in Figure 3), and the density of the background (in code units), respectively. See Figure 3 and text for further details. We have plotted only a section of the cubic box in order to make the particle trajectory clearer.
Figure 2 depicts an example of a particle trajectory in the original MHD turbulence simulation. A characteristic behavior observed in every model is long flights interrupted by successive bounces, where the particle interacts with different magnetic mirrors (Levy-flght-like propagation). As discussed in LX21, only particles with pitch angles above a critical value should interact with magnetic mirrors present along their path (Equation 4). When these conditions are met, the particle reverses its moving direction and bounces between mirrors. In the meantime, the particle also undergoes superdiffusion in the direction perpendicular to the mean magnetic field as a result of the turbulent reconnection of the magnetic fields. Because of the perpendicular superdiffusion, the particle does not remain trapped by mirrors, but diffusively move along the magnetic field when interacting with different mirrors. When the pitch angle becomes sufficiently small due to the pitch-angle scattering, the particle can escape from mirrors and undergo the scattering diffusion. With the stochastic change of pitch angle due to the gyroresonant scattering, the particle stochastically experiences both the mirror diffusion and scattering diffusion.
Figure 3: The first panel depicts the time evolution of the cosine of the pitch angle ($\mu$) for the same particle shown in Figure 2, for complete background magnetic field distribution. The gradient color scheme depicts the portion of the particle trajectory as in Figure 2. Gray line shows the complete variation of $\mu$ over time and the blue line shows the same, but without the fastest variations of $\mu$. The middle panel shows the displacement of the particle parallel to the magnetic field with regard to its initial position, and the bottom panel shows $\zeta/\zeta_{0}$ (Equation 3) over time. Red circles stand for the identified bounces. Black triangles are points that are not considered as bounces due to mirror diffusion (see text for details). Time is normalized by the initial value of the girofrequency $\Omega=c/R_{L}$ (Equation 2), being $c=1$ in the code units.
Figure 4: The same as Figure 3, but with smoothed magnetic field as background.
In Figure 3, we show the time evolution of $\mu$ (top panel), the distance traveled parallel to the magnetic field versus time (middle panel), and $\zeta/\zeta_{0}$ versus time, where $\zeta$ is the adiabatic invariant of the particle, which is conserved when the particle is bouncing in a mirror:
$\zeta(t)=\gamma mv^{2}\frac{1-\mu(t)^{2}}{2|\textbf{B}(t)|},$ (3)
where $\zeta_{0}=\zeta(t=0)$ for the same particle shown in Figure 2.
As an example, Figure 2 shows just a part of the particle’s trajectory. The respective time interval is shown in Figure 3 with the same color coding. Considering this time interval, the second panel of Figure 3 shows that the particle travels upward until it interacts with a mirror, staying more or less in the same height until it is able to continue its trajectory downwards after several consecutive mirroring events. In Figure 2 we see that during these several consecutive bounces, the particle continues to move perpendicular to the magnetic field, interacting with different mirrors. In the lower panel of Figure 3, the respective time interval also shows a nearly constant adiabatic invariant, which reinforces the fact that these are mirroring events.
The peaks right before each bounce point (i.e., when $\mu=0$) in Figure 3, are identified as the beginning of a bounce $\mu_{b}$ is the value of $\mu$ for each of these peaks. $\mu_{b}$ reflects the angle with which the particle starts to interact with a magnetic mirror. In other words, if $\mu_{b}>\mu_{c}$ the particle should not interact with the mirror, where $\mu_{c}=\cos{\alpha_{pitch,c}}$ is the critical value of the pitch angle cossine below which particle mirror diffusion occurs, as defined in Section 1.
For fluctuations induced by fast modes of MHD turbulence, LX21 obtained the critical value of $\mu_{c}$:
$\mu_{c}$ $=$ $\Bigg{[}\frac{14}{\pi}\frac{\delta B_{f}^{2}}{B_{0}}\Big{(}\frac{v}{L\Omega}\Big{)}^{1/2}\Bigg{]}^{2/11},$ (4)
where $\delta B_{f}$ is the perturbed magnetic field of fast modes, $B_{0}$ is the mean magnetic field, $v$ is the particle velocity, $L$ is the turbulence driving scale, and $\Omega$ is the particle’s gyrofrequency.
On the other hand, in the limit of incompressible turbulence $\mu_{c}$ will be dominated by fluctuations caused by pseudo-Alfvén modes:
$\mu_{c,inc}$ $=$ $\sqrt{\frac{\delta B_{s}}{B_{0}+\delta B_{s}}},$ (5)
where $\delta B_{s}$ is the perturbed magnetic field of slow modes. A thorough discussion of all potential scenarios regarding MHD modes dominance over the bounce diffusion process is beyond the scope of this work, but see Section 5.2 of LX21 for a more detailed analysis.
To identify the peaks in $\mu$, we first filter the values of $\mu$ along time using a Butterworth digital filter (Virtanen et al., 2020), which helps us to follow the average behavior of $\mu$ along time. Note that there is a fast variation of $\mu$and that not all bounce points correspond to $\zeta/\zeta_{0}$ constant. This variation is reduced comparing the first panels of Figures 3 and 4. Figure 4 shows the information presented in Figure 3, but for a particle propagating in a smoothed background distribution of the magnetic field. In Figure 3, top, the gray line (original data) has a component that partially hides the average behavior of the particle which is depicted by the blue line (filtered data). In Figure 4, since the magnetic field background is smoothed and scattering is already reduced, the peaks are more easily identifiable.
In turbulence particle trajectories are distorted by scattering. To distinguish such events from mirror diffusion, we introduce the following criterium. Let us recall that $\mu_{b}$ is the peak value of $\mu$ right before a bounce. We can thus evaluate the respective adiabatic invariant of $\mu_{b}$, $\zeta$ using Equation 3. This is shown in the bottom panel of Figures 3 and 4. For a particle with $n$ occurrences of $\mu=0$, and each peak occurring at times $t_{1}<t_{2}<...<t_{i}<...<t_{n}$, we compare $\zeta(t_{i})/\zeta_{0}$ with $\zeta(t_{i-1})/\zeta_{0}$ and $\zeta(t_{i+1})/\zeta_{0}$ (to simplify notation, consider $\zeta_{i}=\zeta(t_{i})/\zeta_{0}$). With this, for a given bounce at time $t_{i}$, $\mu_{b}(t_{i})$ is discarded if $|\zeta_{i}-\zeta_{i-1}|>0.2$ and $|\zeta_{i}-\zeta_{i+1}|>0.2$. In other words, if the variation of $\zeta_{i}$ is bigger than $0.2$ compared to its neighbors, the bounce is discarded [^2] The value of 0.2 was arbitrarily chosen to ensure that we had a maximum of $\sim 20\%$ variation in the adiabatic invariant for each identified $\mu_{b}$.. For the first identified point, $\zeta_{1}$, we compare with the values of the next two identified points whilst for the last one, $\zeta_{n}$, we compare with the two before. If there are less than 3 bounce events, we simply accept the values of identified $\mu_{b}$.
3.2 Pitch angle distribution of $\mu_{b}$
To evaluate the pitch angle distribution of bounce events, we will use the values obtained of $\mu_{b}$ and compare with the maximum critical value (LX21):
$\mu_{c,max}\simeq\sqrt{\frac{\delta B}{\delta B+B_{0}}},$ (6)
where $\delta B$ is the perturbed field locally defined by the position of the particle and $B_{0}$ is the mean magnetic field [^3] We note that this equation corresponds to the maximum value of $\mu_{c}$ which is attained when $\mu_{c}$ becomes independent of the particle energy (LX21)..
Considering what was discussed in Section 3.1 and the information in Figures 3 and 4, we can produce the distribution of $\mu_{b}$ and its probability density distributions (PDFs) for all models.
Figure 5: Comparison between the expected values of $\mu_{c,max}$ (Equation 6) and the identified bounce values from the simulations $\mu_{b}$. The orange line indicates where $\mu_{c,max}=\mu_{b}$. The color map shows the density of points in the diagrams. Top panels show the results for models with the complete magnetic field distribution, while the bottom panels show the results for models with the smoothed distribution of magnetic fields. On the left hand side we show models with $\mu_{0}=0.2$, while on the right hand side we show models with $\mu_{0}=0.8$.
Figure 6: Probability density distributions of $\mu_{b}$ (see text for further details). Colors indicate different Larmor radii of the particles. The top panels show the PDFs for models with the complete background magnetic field, whilst the bottom panels show PDFs for models with the smoothed magnetic field as background. On the left hand side we show models with $\mu_{0}=0.2$, while on the right hand side we show models with $\mu_{0}=0.8$.
Figure 5 compares values of $\mu_{b}$ with the respective $\mu_{c,max}$ for every bounce identified in the models with $R_{L}=0.03L_{0}$. The top panels show the results for models with complete distribution of magnetic fields, while the bottom panels show the same for models that consider the smoothed distribution of magnetic fields. When $\mu_{0}$ is small, the number of bounces is generally higher than when $\mu_{0}$ is closer to 1. Comparing the upper and bottom panels, it is clear that the lack of scattering plays an important role in slowing down the diffusion parallel to the magnetic field, as evidenced by the models with a smoothed background.
In the smoothed models in Figure 5 (panels (c) and (d)) we see that the identified $\mu_{b}$ events are mostly smaller than the corresponding predicted maximum values $\mu_{c,max}$. This means that the particles are bouncing from larger amplitude mirrors in the MHD cascade. As shown in Figure 4, the peaks maintain similar values while the particle bounces through successive mirrors. Also, $\zeta$ stays almost constant during this process, which indicates that $\mu_{b}$ is a good proxy for the bouncing process. The probability of a bounce occurring for a certain value of $\mu$ is not very sensitive to the initial conditions, unless scattering is suppressed. As evidenced by panels (c) and (d) of Figure 5, where we have a smoothed distribution of the magnetic field, particles interact less with the mirrors if the initial $\mu$ is large ($\mu_{0}=0.8$).
This scenario is quite different when scattering is more important (panels (a) and (b) of Figure 5). The initial pitch angle is less important for the overall behavior of the test particles and the identification of bounces with constant $\zeta$ is harder due to the presence of variations that hide the average behavior of the particle, introduced by pitch angle scattering. Comparing $\mu_{b}$ with its respective $\mu_{c,max}$ for this case shows a larger number of points below the orange line in Figure 5. Points identified here as bounce have already been compared with their neighbors, according to what was discussed in Section 3.1. These deviations can generally be explained by one of two reasons: either the particles are scattered and change directions or, due to scattering, $\mu$ later becomes smaller than $\mu_{c,max}$ along the particle trajectory and then the particle bounces while interacting with a mirror. In the former case, the misidentifying occurs due to our criterion not being strict enough. Still, the number of mirroring events are way more numerous than the events of bounces caused by scattering. In the latter case, we are not able to track the beginning of a bounce by looking at the peaks of $\mu$ versus time, so that defining a $\mu_{b}$ for these cases can be somewhat arbitrary. Still, these events are not frequent enough to interfere in our analysis.
Figure 6 shows the PDFs of $\mu_{b}$ for all models. We calculate the PDF for all bounces integrated in time, after $t\times\Omega=2000$. Different colors represent different initial Larmor radii for the particles, with panels on the left-hand side showing models with an initial pitch angle such that $\mu_{0}=0.2$ and on the right-hand side with $\mu_{0}=0.8$. The top two panels show the results for models with the complete magnetic field whilst the bottom panels show results for models with the smoothed field as the background.
The distributions show that all models have a drop in probability for $\mu_{b}\gtrsim 0.6$. Also, most models show a peak in the distribution between $0.2$ and $0.3$.
For the smoothed model r03mu80k5, only $\sim 50\%$ of the particles interact with a magnetic mirror, and about $1/5$ of the bouncing particles suffer only one or two bounces along the trajectory [^4] We note that in this case due to the lack of at least three bouncing events, we classified these events as mirroring. This introduces two points that need to be considered. First, the fraction of mirroring particles in this model is much smaller than in other models, where the pitch angle of the particle either is smaller at the beginning or resonant scattering changes the particle’s pitch angle enough that their probability of interacting with a mirror is higher. Second, as the particles travel inside the domain with higher values of $\mu$, they will see changes in the magnetic field and will only interact when the conditions for bounce are met, which will be skewed towards higher $\mu$ values due to the initial condition. The probability of a mirroring event occurring with $\mu>0.6$ is still small, but there are further complications that needs to be considered as we will discuss in the next section.
3.3 Particle diffusion in MHD turbulence
Figure 7: Top panels show the average squared displacement as a function of time normalized by the giro-frequency of the particles around the mean magnetic field. The second row shows the time derivative of the top panels, $\Delta\tilde{z}^{2}/\Delta t=\Delta<(z-z_{0})^{2}>/\Delta t$, giving the parallel diffusion coefficient when it becomes constant. The third row shows the average value of $\mu$ over all particles. The fourth and last row shows the average value of $\zeta/\zeta_{0}$ (Equation 3). Different colors represent different particles Larmor radii according to the legend. Time in x-axis is normalized by the initial value of the girofrequency $\Omega=c/R_{L}$ (eq. 2), with $c=1$ in code units.
Figure 8: Same as Figure 7, but for models that consider the smoothed distribution of the background magnetic field. Different colors represent different Larmor radii according to legend. We note that for small Larmor radius, diffusion in $\mu$ occurs slowly, leading to a delay in reaching the steady state within the specified time (see text for more details).
From top to bottom panels, Figure 7 compares the squared displacement of particles, the second panel shows the evolution of first derivative of the upper panel, $\Delta\tilde{z}^{2}/\Delta t$ (where we write $\tilde{z}^{2}=<(z-z_{0})^{2}>$ to concise the notation), the average value of $\mu$ over all particles, and the average value of $\zeta/\zeta_{0}$ (Equation 3) versus the normalized time $(t\times\Omega)$. The colors represent models with $R_{L}/L_{0}=0.03,0.06$ and $0.10$ for blue, orange and green, respectively. Note also that $\Omega$ is different for each Larmor radius. All models in this figure consider the complete magnetic field distribution of the background.
The changes in $\Delta\tilde{z}^{2}/\Delta t$ over time in Figure 7 highlight the influence of the initial pitch angle. Initially, particles with low $\mu_{0}$ exhibit a smaller diffusion coefficient, with models starting with smaller Larmor radii requiring slightly more time to reach $<\mu>\simeq 0.5$. However, all the models incorporating the full magnetic field distribution demonstrate a stable diffusion process around $t\times\Omega\sim 10^{2}$. Note that when $\tilde{z}^{2}=<(z-z_{0})^{2}>$ increases at slower rate, the associated diffusion coefficient $D_{\parallel}=\Delta\tilde{z}^{2}/\Delta t$ becomes close to $\sim 10^{-2}$ for all models. However, there is actually a mild dependence of this saturation value with the Larmor radius of the model. This dependence is clearer in Figure 9 (bottom panel, Section 4), which does not use a logarithmic scale for the plot.
Figure 8 depicts the same information as Figure 7, but for models featuring a smoothed distribution of the magnetic field as the background. The effects described earlier are more pronounced in these models. Specifically for $R_{L}/L_{0}=0.03$, there is a significant impact from magnetic mirrors on particles with $\mu_{0}=0.2$ as the influence of pitch-angle scattering diminishes. This is reflected in the time required for $\Delta\tilde{z}^{2}/\Delta t$ to stabilize. Towards the end of the observed period, $\Delta\tilde{z}^{2}/\Delta t$ is approximately one order of magnitude smaller for $\mu_{0}=0.2$ compared to $\mu_{0}=0.8$.
The average behaviour of the adiabatic invariant in Figure 8 confirms that particles with smaller $\mu_{0}$ suffer mirroring for a longer duration initially. The average value of $<\zeta/\zeta_{0}>$ for models with $\mu_{0}=0.2$ begins to exhibit more significant variations only after $t\times\Omega=10^{2}$, whereas for models with $\mu_{0}=0.8$, variations commence after $t\times\Omega=10^{1}$.
This behavior can be elucidated by considering the following. Particles do not interact with fluctuations smaller than their Larmor radius, and pitch angle scattering is associated with interactions with fluctuations having a wavelength $\lambda\sim R_{L}$. At the same time, large-scale perturbations induce the formation of magnetic mirrors. Referring back to Figure 1, $R_{L}/L_{0}=0.03$ corresponds to $k\sim 30$, which is indicative of magnetic energy smaller than $10^{-13}c.u.$. In essence, for $\mu_{0}=0.8$, particles can travel along magnetic field lines at much higher velocities, barely encountering mirrors initially due to their small pitch angle. However, as they traverse through small fluctuations, whether originating from numerical noise or residuals from smoothing, they experience scattering, initiating variations in $<\zeta/\zeta_{0}>$ earlier compared to the scenario when $\mu_{0}=0.2$. In the latter case, particles remain confined by a series of magnetic mirrors, predominantly diffusing perpendicular to the magnetic field.
The effect of mirrors and scattering over the particles with smaller $R_{L}$ in Figures 7 and 8 is two-fold, as evidenced by the different magnetic field backgrounds. Mirrors can be very effective in confining particles in a system where scattering is suppressed at the same time that the particles can escape quickly in these same systems if they start their propagation in a direction close to parallel to the magnetic field. Scattering serves as an regulating mechanism, facilitating mirror diffusion along the magnetic field.
An illustrative demonstration of this regulatory effect can be observed in the models depicted in Figure 8, with $R_{L}/L_{0}$ values of 0.06 and 0.10 (corresponding to $k\sim 17$ and $k\sim 10$, respectively). The temporal evolution of $\Delta\tilde{z}^{2}/\Delta t$ in these models closely resembles that of the scenario with a complete background magnetic field distribution (Figure 7), albeit with some nuanced distinctions. As the Larmor radius increases, the influence of the smoothing of the magnetic field distribution at small scales becomes less effective. Models with $R_{L}/L_{0}=0.06$ experience less scattering compared to those with $R_{L}/L_{0}=0.10$, yet the former ends up with a slightly higher diffusion coefficient than the latter. This seems to suggest that scattering becomes more efficient at higher energies, but we need to take into account that the dissipative scale for this background is around $\sim 0.2L_{0}$ (see Section 2.2). As the Larmor radius of the models becomes close to the dissipative scale, they can be scattered more efficiently even at higher $\mu$. This can also explain why the green models ($R_{L}/L_{0}=0.10$) seem to be better confined than the orange ones ($R_{L}/L_{0}=0.06$). At a small $\mu_{0}$, the non-resonant mirroring can well confine particles even when the Larmor radius is much smaller than the dissipative scale (see also Figure 10).
Finally, Figures 7 and 8 also indicate that, although the particles can achieve a close to constant $\Delta\tilde{z}^{2}/\Delta t$ very quickly, which is indicative of a stable diffusion process, the entire process actually occurs over 5 to 10 times the size of the simulated box. This means that this initial phase of the diffusion, which can be heavily affected by mirrors, can occur over a considerable sized region around the source. Note however that this depends on how efficient resonant scattering is and it will also depend on the particle’s energy. Applications of these results and comparisons to observations will be discussed in the next section.
4 Discussion
$B(\mu G)$ $L_{0}(pc)$ $R/L_{0}$ E(PeV) $D_{\parallel}$ $D_{\parallel,phys}(cm^{2}/s)$
3.2 1 0.03 0.09 0.005 $4.63\times 10^{26}$
3.2 1 0.06 0.18 0.006 $5.55\times 10^{26}$
3.2 1 0.10 0.30 0.008 $7.40\times 10^{26}$
3.2 5 0.03 0.44 0.005 $2.31\times 10^{27}$
3.2 5 0.06 0.89 0.006 $2.78\times 10^{27}$
3.2 5 0.10 1.48 0.008 $3.70\times 10^{27}$
10 5 0.03 1.40 0.005 $2.31\times 10^{27}$
10 5 0.06 2.80 0.006 $2.78\times 10^{27}$
10 5 0.10 4.67 0.008 $3.70\times 10^{27}$
Table 2: Diffusion coefficients as obtained from our models with complete magnetic field distribution (Figures 7 and 9), for different system sizes and magnetic fields, and different particle’s energy.
Figure 9: The same as in Figure 7 for a complete magnetic field distribution, but considering the physical values of the particles displacement and the parallel diffusion coefficient, taking a system with size $L_{0}=3pc$ (see Equation 7).
Figure 10: The same as in Figure 9, but considering a smoothed background distribution of the magnetic field (as in Figure 8).
In this work we have explored the effects of mirror diffusion of CRs by combining 3D MHD simulations of a supersonic, sub-Alfvenic turbulent molecular cloud a with test particle simulations, considering initial pitch angles corresponding to $\mu=0.2$ and $0.8$.
Particles injected into the turbulent domain undergo both bouncing and scattering processes. We have observed that mirror diffusion significantly influences their parallel diffusion component relative to the mean magnetic field, validating the LX21 predictions.
Analyzing their $\mu_{b}$, defined as the last maximum value of $\mu$ where their adiabatic invariant $\zeta$ exhibits minimal variation over time, reveals that the majority of bounces occur for $\mu_{b}<0.6$, however this should depend on the background magnetic field.
The initial diffusion of CRs is influenced by their initial pitch angle, with those starting closer to $90^{\circ}$ (or $\mu_{0}$ closer to 0) being more prone to mirror diffusion and requiring longer diffusion times, while those with smaller pitch angles (larger $\mu_{0}$) can escape the system more rapidly. This disparity also affects the time required for particles to attain normal diffusion. Particles beginning with pitch angles near $90^{\circ}$ achieve a nearly constant $\Delta\tilde{z}^{2}/\Delta t$ at a slower rate (see Figures 7 and 8). Suppression of pitch angle scattering can further accentuate the role of initial pitch angle in defining the particle diffusion regime, depending on its Larmor radius.
We can convert the parallel diffusion coefficient $D_{\parallel}$ (defined as the value of $\Delta\tilde{z}^{2}/\Delta t$ after it stabilizes) obtained from the simulations in code units into physical units, by considering an arbitrary physical size for the system, as follows:
We can convert the parallel diffusion coefficient $D_{\parallel}$ (the value of $\Delta\tilde{z}^{2}/\Delta t$ after it achieves a nearly stable value), obtained from the simulations in code units, into physical units, considering an arbitrary physical size for the system, as follows:
$\begin{split}D_{\parallel,phys}[cm^{2}/s]&=D_{\parallel}\cdot L_{0}\,c\\ &=D_{\parallel}\cdot 2.776\times 10^{29}\,\Big{(}\frac{L_{0}}{3pc}\Big{)},\end{split}$ (7)
where $L_{0}$ denotes the physical dimension of the system under examination, and $c$ is the light speed.
We can also convert the magnetic field from our simulations in code units into physical units, as follows:
$\begin{split}B[\mu G]=3.215\,\Big{(}\frac{n}{10cm^{-3}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{c_{s}}{1.9\cdot 10^{4}cm/s}\Big{)}\\ \Big{(}\frac{M_{s}}{7.0}\Big{)}\Big{(}\frac{M_{A}}{0.6}\Big{)}^{-1},\end{split}$ (8)
where $n$ is the number density, and $c_{s}$ is the sound speed of the system. The values adopted in this equation are representative of a molecular cloud region, but in Table 2 we also consider parameters from other systems.
Considering that particles travel close to speed of light, we can write:
$R_{L}[pc]$ $=$ $1.084\frac{(E/1PeV)}{(B/1\mu G)}\Big{(}\frac{v_{\perp}}{c}\Big{)},$ (9)
$\Omega[s^{-1}]$$\approx$ $9.743\times 10^{-9}\Big{(}\frac{1pc}{R_{L}}\Big{)},$ (10)
where $E$ is the energy of the particle, $\Omega$ its gyrofrequence, and $R_{L}$ its Larmor radius in second and pc units, respectively. As stressed before, we have assumed $v_{\perp}\approx c$.
As emphasized in Section 1, recent findings from VHE gamma-ray observatories like H.E.S.S., HAWC and LHAASO have revealed extensive diffuse emissions reaching a few to hundreds of TeV and even PeV within our galaxy (Abeysekara et al., 2017; Huang et al., 2018; Abeysekara et al., 2021; Cao et al., 2021), without definitive identification of their sources. Consequently, comprehending CR acceleration and diffusion in the surrounds of embedded sources has become critically important for interpreting the observations. Traditional models and assumptions to characterize CR diffusion are insufficient to elucidate these emissions.
To compare our results with observations, we present in Figure 9 the same diagrams as in Figure 7 for the complete magnetic field distribution, but depicting the physical parallel diffusion coefficient, $(\Delta\tilde{z}^{2}/\Delta t)|_{phys}$ (see Equation 7), as a function of the time ($L_{0}/c$) in seconds. For the physical size of the system, we have adopted $L_{0}=3pc$ in this figure.
Combining the results of Figures 7 and 9, with Equation 9 of the Larmor radius to evaluate the particle’s energy for a given background magnetic field, we can obtain the corresponding parallel diffusion coefficient for the particle, as predicted from our models. Table 2 shows $D_{\parallel,phys}$ for particles with different energies and magnetic fields in systems with different sizes. We also observe that particles travel approximately 10 times the scale $L_{0}$ before attaining the saturation point (i.e., a nearly constant value) of the parallel diffusion coefficient. Throughout the initial stage, the diffusion of the particles varies based on their initial pitch angle, with greater susceptibility to mirror effects when the pitch angle approaches $90^{\circ}$.
We note that $D_{\parallel,phys}$ is consistent with the implied values to explain VHE observations of extended sources like Geminga and PSR B0656+14 (Abeysekara et al., 2017). Their combined halo has an inferred diffusion coefficient of $4.5\times 10^{27}cm^{2}/s$ for 100 TeV CRs within an extended zone with tens of parsec (Abeysekara et al., 2017). Similarly, Huang et al. (2018) propose a diffusion coefficient $\lesssim 10^{28}\,cm^{2}/s$ for electrons of $\sim 10$TeV in a region of a few tens of parsecs around Vela X.
Interestingly, our results from the models featuring smoothed background magnetic field distribution (Figures 8 and 10) could also offer insights into systems like the nuclear regions of starburst galaxies. For example, Krumholz et al. (2020) argues that, as in molecular clouds within our Galaxy, these systems are likely to possess sufficiently low ionization fractions. Consequently, ion-neutral damping would halt turbulent cascades at scales much larger than the gyroradius of CRs with energies in the range of a few hundred TeV. Still, CRs could scatter off turbulence self-generated by streaming instability. The competition between the streaming instability and ion-neutral damping leads to a CR streaming speed equal to the ion Alfvén speed, which is independent of CR energy up to $\sim 1TeV$. Diffusion should occur only due to field line random walk along the magnetic field, yielding a diffusion coefficient of approximately $10^{27}-10^{28}\,\text{cm}^{2}/\text{s}$, in this energy range. While the behavior of these CRs would align with the results shown in Figures 8 and 10 in the absence of streaming instability feedback, proper modeling of diffusion in such environments necessitates the incorporation of CR feedback effects too.
Considering all the preceding discussions, it’s worth noting that while pitch angle scattering alone fails to generate sufficiently small diffusion coefficients to align with observations, mirror diffusion alone also falls short. However, the interplay between both mechanisms appears to yield the slowest diffusion coefficients. This is evident when examining, for instance, all models with $R_{L}/L_{0}=0.03$ (indicated by blue curves) in Figures 9 and 10. In the former, $D_{\parallel}$ stabilizes around $\sim 10^{27}\,\text{cm}^{2}/\text{s}$, whereas in the latter, it continues to increase, without yet displaying indications of approaching a nearly constant value by the end of the considered timeframe.
The variation of $\mu$ due to resonance scattering allows for more particles to interact with mirrors, resulting in a net coefficient that is smaller than the one obtained from each process separately. Therefore, mirror diffusion needs to be taken into account with pitch angle scattering to properly model CR diffusion.
Zhang & Xu (2023) have recently focused on the mirror diffusion alone, testing on the pitch-angle-dependent mean free path of mirror diffusion. Here we have studied the overall propagation behavior with both mirror and scattering diffusion taken into account. We find that the propagation behavior in the early time, e.g., in the vicinity of CR sources, depends on the initial pitch angle distribution. Away from CR sources, it transits from the initial parallel superdiffusion to the parallel normal diffusion.
Lastly, it is worth noting that recent studies by Lemoine (2023) and Kempski et al. (2023) have examined the implications of intermittent bends and reversals in magnetic fields on CR diffusion. They conclude that low-energy particles experience more efficient confinement than high-energy particles, attributed to the absence of pitch angle scattering and isotropization. While employing different configurations for the background magnetic field — such as utilizing incompressible turbulence in the former or omitting a guide field in the latter — their findings exhibit similarities to the outcomes observed in our models, especially in the ones with smoothed background distribution, particularly regarding $R_{L}/L_{0}=0.03$. The absence of pitch angle scattering is a significant factor influencing fast particle escape rates ($\mu_{0}=0.8$) or prolonged mirroring ($\mu_{0}=0.2$). However, further investigation is warranted to explore the intersections between their studies and ours.
5 Conclusions
In this work, we have numerically demonstrated the process of mirror diffusion theoretically predicted in LX21. This effect is particularly significant for particles with large pitch angles. We found that in combination with resonance scattering, the particle transport is akin to a Levy-flight-like propagation. Our numerical results agree with those in Zhang & Xu (2023), but extended them accounting for both mirror and scattering diffusion to assess their combined effects aiming at applications to realistic astrophysical conditions.
We have found that when pitch angle scattering is minimized, the impact of mirror diffusion depends on the particle’s initial pitch angle, for a significant amount of time. This effect is especially evident for lower energy particles.
Furthermore, our findings have revealed that the range of pitch angles subjective to mirroring aligns with the predictions derived from the models of Xu & Lazarian (2020) and LX21, and the estimated physical values of the diffusion coefficient, $\lesssim 10^{27}cm^{2}/s$, fall within recent constraints imposed by VHE energy observations.
The combination of resonance scattering with mirror diffusion leads to a slower diffusion compared with the case with scattering alone of particles and we contend that mirror diffusion should be considered when studying CR propagation within turbulent magnetized environments. Moving forward, we intend to integrate these effects into comprehensive simulations of CR propagation in massive star-forming regions in order to quantify their VHE emission.
L.B-M. acknowledges support from the Brazilian funding agency CAPES, CNPq and also the kind hospitality of Professor Alex Lazarian and the University of Wisconsin - Madison; E.M.G.D.P. acknowledges support from the Brazilian funding agencies FAPESP (grants 2013/10559-5 and 2021/02120-0) and CNPq (grant 308643/2017-8); This research was also supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). The simulations presented in this work were performed in the cluster of the Group of Plasmas and High-Energy Astrophysics of IAG-USP (GAPAE) acquired with support from FAPESP (grant 2013/10559-5).
Data Availability
The data of this article will be available upon request to the corresponding author.